# Nonlinear Kinematics of Recursive Origami Inspired by Spidron

### v1

@@ 5 Conclusion
The isotropic rigid folding of the Spidron that converges to the finitely folded state could be used to store sheets in the compact configuration as the well-known flasher origami, which is not rigidly foldable without cuts [Lang et al. 16]. The quick convergence to a finitely folded state implies that the structure folds to a desired shape automatically by just actuating the outermost boundary of the developed structure. In addition, we observed that a real model of the Spidron can have multi-stability as the hypar origami [Filipov and Redoutey 18, Liu et al. 19], which is caused by the mode-switching between pro- and anti-rotation modes. Because mode-switching results in the switching of the dominating dynamical system, it can switch the mechanical behavior of the structure. From these engineering perspectives, exploring the mechanical properties of our pattern is one of the possible future directions. Another direction would be the generalization of the boundary shapes to such as rectangular [Demaine et al. 99], hexagonal boundaries like the original Spidron [Chen et al. 23], or circular boundaries. This generalization would bring new mathematical/mechanical properties.

### v2

@@ 5 Conclusion
The isotropic rigid folding of the Spidron that converges to the finitely folded state could be used to store sheets in the compact configuration as the well-known flasher origami, which is not rigidly foldable without cuts or the triangulation [Lang et al. 16]. The quick convergence to a finitely folded state implies that the structure folds to a desired shape automatically by just actuating the outermost boundary of the developed structure. In addition, we observed that a real model of the Spidron can have multi-stability as the hypar origami [Filipov and Redoutey 18, Liu et al. 19], which is caused by the mode-switching between pro- and anti-rotation modes. Because mode-switching results in the switching of the dominating dynamical system, it can switch the mechanical behavior of the structure. From these engineering perspectives, exploring the mechanical properties of our pattern is one of the possible future directions. Another direction would be the generalization of the boundary shapes to such as rectangular [Demaine et al. 99], hexagonal boundaries [Chen et al. 23] like the original Spidron, or circular boundaries. This generalization would bring new mathematical/mechanical properties.